This is the question: Science and medicine. A bicyclist rode into the country for 5 h. In returning, her speed was 5 mi/h faster and the trip took 4 h. What was her speed each way?
For some reason and I am stumped. I don't know if I'm not reading it correctly, or if I'm missing information.
Can you help?
Thanx, Gerri

Answer by checkley71(8403)

DISTANCE=RATE*TIME
SINCE THESE DISTANCES ARE EQUAL THEN:
X*5=(X+5)4
5X=4X+20
5X-4X=20
X=20 MPH FOR THE 5HOUR TRIP INTO THE COUNTRY.
20+5=25 MPH FOR THE RETURN TRIP.
PROOF
20*5=(20+5)*4
100=25*4
100=100